OpenStudy (anonymous):

What is Andrea's rate in mi/h? Andrea can ride her bike 2 mi/h faster than Ruby. It takes Andrea 48 min less to travel 50 mi than it does Ruby. What is Andrea's rate in mi/h? explain steps :D

jimthompson5910 (jim_thompson5910):

let x = Ruby's speed, y = Ruby's time Ruby: d = rt 50 = x*y y = 50/x Andrea: d = rt 50 = (x+2)*(y-48) --------------------- Now plug in y = 50/x into the last equation above and solve for x